OpenStudy (anonymous):

What is the approximate mass of an oxygen atom in atomic mass units? What is the approximate mass of two oxygen atoms? How about an oxygen molecule? @Abhisar

OpenStudy (abhisar):

No of protons in an Oxygen atom = 8, No. of neutrons in an oxygen atom = 8 Atomic mass = 8+8=16u

OpenStudy (abhisar):

u=symbol for atomic mass unit

OpenStudy (anonymous):

oxygen is 15.999? 15.999*2=31.998?

OpenStudy (abhisar):

yes, mass of one oxygen molecule will be 16*2 because one molecule of oxygen contains 2 atoms

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

no it was a guess

OpenStudy (abhisar):

Mass of one O = 16 Mass of 2 O = 32 Mass of one Oxygen molecule (O\(_2)\) = 32

OpenStudy (abhisar):

All clear ?

OpenStudy (anonymous):

so a molecule contains 2?

OpenStudy (abhisar):

Yes, one oxygen molecule contain 2 atom, but one nitrogen molecule 3 atom.

OpenStudy (anonymous):

ok got it

OpenStudy (abhisar):

it depends upon the molecular formula

OpenStudy (anonymous):

ok